Sorry to submit the same question but I think I may have typed it incorrectly. The question relates to Newton's Synthesis. The equation for the force of gravity is equal to the equation for centripetal force. I am asked to solve for "G" given:

G*(Mm/r^2)= mv^2/r


Thank you!

Answer by Boreal(15235)

This is GMm/r^2=mv^2/r
The m and one r cancel
so GM/r=v^2
G=v^2*r/M